Session: Feb 2006 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Medicina (Fiocchi Franco at Mon Feb 27 03:48:36 2006) Success :

Onsala (Michael Lindqvist at Mon Feb 27 05:56:55 2006) Success :
Bad calibration at 17:40 UT, could have been RFI. Otherwise no known problems.

Torun (Kaz Borkowski at Mon Feb 27 06:29:53 2006) Success :
No known problems.

Westerbork (Tony Foley at Mon Feb 27 12:51:50 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems

Effelsberg (A. Polatidis at Tue Feb 28 10:33:27 2006) Minor failures : WIND
Stopped 1615-1745 UT due to snow

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:17:40 2006) Failure : RECO ANT
The MkII telescope was used (Jb2). Offsets were incorrectly set hence sources probably out of telescope beam. Disk pack JOD-0023 subsequently developed "directory flawed or missing" error (during experiment ea035a). No data visible on disk pack but may be recoverable.

Cambridge (Alastair Gunn at Fri Mar 3 17:18:05 2006) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 7 15:58:46 2006) Success :

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)